Sabas Company has 20,000 shares of $100 par, 2% cumulative preferred stock and 100,000 shares of $50 par common stock. The following amounts were distributed as dividends:

Year 1 $10,000
Year 2 45,000
Year 3 90,000


Determine the dividends per share for preferred and common stock for the first year.

a.$0.50 and $0.10

b.$2.00 and $0

c.$0.50 and $0

d.$0 and $0.10

A corporation has 49,347 shares of $38 par stock outstanding that has a current market value of $350 per share. If the corporation issues a 4-for-1 stock split, the market value of the stock will fall to approximately

a.$312.00

b.$87.50

c.$9.50

d.$1,400.00
Answer #1

Annual Preferred Dividend = 20,000*2%*100 = 40,000

Dividend Paid

to Preferred Stock =$10,000

to Preferred Stock = 0

Dividend Per share

Preferred = 10,000 /20,000 = $0.50

Common Stock = 0

Answer = C) $0.50and $0

Stock Split 4 to 1

Market Valur after STock Split = Current Market Price / Shares = 350, /4 = $87.50

Answer = B) $87.50